# Towards Open Vocabulary Learning: A Survey

### v1

@@ 2.3 Tasks, DataSets, and Metrics
$\bullet$ Dataset. For object detection, the common datasets are COCO [5] and LVIS [40]. Recently, a more challenging dataset, v3Det [41] with more than 10,000 categories is proposed. For image segmentation, the most commonly used datasets are COCO [5], ADE20k [42], PASCAL-VOC 2012 [43], PASCAL-Context [44], and Cityscapes [45]. For video segmentation and tracking, the frequently used datasets are VSPW [46], Youtube-VIS [47], LV-VIS [48], MOSE [49], and TAO [50].
$\bullet$ Metrics. For detection tasks, the commonly used metrics are mean average precision (mAP) and mean average recall (mAP) for both base classes and novel classes. As for segmentation tasks, the commonly used metrics are mean intersection over union (mIoU) for semantic segmentation, mask-based mAP for instance segmentation, and panoptic quality (PQ) for panoptic segmentation.
@@ 2.4 Related Research Domains
$\bullet$ Long-tail Object Detection and Instance Segmentation. The objective of this task is to address the challenge of class imbalance in instance segmentation. Many approaches tackle this issue through techniques such as data re-sampling [40, 51, 52], loss re-weighting [53, 54, 55, 56, 57], and decoupled training [58, 59]. Specifically, some studies [51] employ image-level re-sampling. However, these methods tend to introduce bias in instance co-occurrence. To address this issue, other works [60, 52, 61] focus on more refined re-sampling techniques at the instance or feature level. Regarding loss re-weighting, most studies [53, 54, 62] rebalance the ratio of positive and negative samples during training. Additionally, decoupled training methods [58, 59] introduce different calibration frameworks to enhance classification results. Resolving long-tail object detection can lead to improved accuracy in detecting novel classes. However, the proposed methods currently cannot be applied to the detection of novel classes.
$\bullet$ Few-Shot Detection and Segmentation: Few-shot object detection aims to expand the detection capabilities of a model using only a few labeled samples. Several approaches [63, 64, 65, 66, 67, 68, 69, 70, 71] have been proposed to advance this field. Notably, TFA [72] introduces a simple two-phase fine-tuning method, while DeFRCN [73] separates the training of RPN features and RoI classification. Moveover, SRR-FSD [74] combines multi-modal inputs while LVC [75] proposes a pipeline to generate additional examples for novel object detection and train a more robust detector. Few-shot segmentation comprises Few-Shot Semantic Segmentation (FSSS) and Few-Shot Instance Segmentation (FSIS). FSSS involves performing pixel-level classification on query images. Previous approaches [76, 77, 78, 79, 80, 81, 82] typically build category prototypes from support images and segment the query image by computing the similarity distance between each prototype and query features. On the other hand, FSIS aims to detect and segment objects with only a few examples. FSIS methods can be categorized into single-branch and dual-branch method. The former [83, 84] primarily focuses on designing the classification head, while the latter [85, 86, 87] introduce an additional support branch to compute class prototypes or re-weighting vectors of support images. This support branch assists the segmenter in identifying target category features through feature aggregation. For example, Meta R-CNN [87] performs channel-wise multiplication on RoI features, and FGN [86] aggregates channel-wise features at three stages: RPN, detection head, and mask head. However, few-shot learning methods still require examples of novel classes during training, and such data or annotations may not be available.
@@
$\bullet$ Zero-Shot Detection and Segmentation: This task aims to segment classes that have not been encountered during training. Two streams of work have emerged: discriminative methods [88, 89, 90, 91, 92] and generative methods [93, 94, 95]. Representative works in this field include SPNet [88] and ZS3Net [93]. SPNet [88] maps each pixel to a semantic word embedding space and projects pixel features onto class probabilities using a fixed semantic word embedding [96, 97] projection matrix. On the other hand, ZS3Net [93] first trains a generative model to produce pixel-wise features for unseen classes based on word embeddings. With these synthetic features, the model can be trained in a supervised manner. Both of these works treat zero-shot detection and segmentation as a pixel-level zero-shot classification problem. However, this formulation is not robust for zero-shot learning, as text embeddings are typically used to describe objects/segments rather than individual pixels. Subsequent works [98, 94, 99, 32, 95, 100] follow this formulation to address different challenges in zero-shot learning. In a weaker assumption where unlabeled pixels from unseen classes are available in the training images, self-training [93, 90] is commonly employed. Despite promising results, self-training often requires model retraining whenever a new class appears. ZSI [101] also employs region-level classification for bounding boxes but focuses on instance segmentation rather than semantic segmentation. More recently, PADing [32] proposes a unified framework to tackle zero-shot semantic segmentation, zero-shot instance segmentation, and zero-shot panoptic segmentation.
@@
Most approaches in open vocabulary learning are based on zero-shot learning settings, such as replacing the fixed classifier with language embeddings. However, these methods struggle to generalize well to novel classes due to the absence of novel class knowledge. Consequently, their performance is limited, and they are not yet practical for real-world applications.
@@ 3 Methods: A Survey
Figure 3: An illustration of a common architecture in open vocabulary object detection and segmentation. The vision model predicts a class embedding for each box/mask. The embeddings are compared to a set of class embeddings generated by a VLM text model like CLIP or ALIGN, using dot products. The class with the highest score is chosen as the predicted label for the object. Note that while humans define the set of possible object classes, during training, the system only has access to a limited set of ”base” classes.
3.1 Concept Definition
Overview. In this section, we first review several basic concepts and network architecture as preliminary knowledge by extending close-set detectors into open vocabulary detectors. Then we sequentially survey the six subsidiary tasks, including object detection  (Sec. 3.3), segmentation (Sec. 3.4), video understanding (Sec. 3.5), 3D scene understanding (Sec. 3.6), and closely related tasks (Sec. 3.7). Note that we only record and compare the most representative works. We list numerous works in Tab. I and Tab. II. Moreover, since there are also several similar tasks (Sec. 3.7), we also survey and compare several tasks, including class agnostic detection and segmentation, open world object detection, and open-set panoptic segmentation.
@@ 3.2 Preliminary
Pixel-based Object Detection and Segmentation. In general, it can be divided into two different aspects: semantic-level tasks and instance-level tasks, where the difference lies in whether to distinguish each instance.
@@
For the former, we take semantic segmentation as an example. It was typically approached as a dense pixel classification problem, as initially proposed by FCN [102]. Then, the following works are all based on the FCN framework [103]. These methods can be divided into the following categories, including better encoder-decoder frameworks [104, 105], larger kernels [106, 107], multiscale pooling [108, 109], multiscale feature fusion [110, 111, 112, 113, 114], non-local modeling [115], and better boundary delineation [116, 117, 118]. After the transformer was proposed, with the goal of global context modeling, several works propose the variants of self-attention operators to replace the CNN prediction heads [119, 120].
For the latter, we take object detection and instance segmentation for illustration. Object detection aims to detect each instance box and classify each instance. It mainly has two different categories: two-stage approaches and one-stage approaches.
@@
Two-stage approaches [121, 2] rely on an extra region proposal network (RPN) to recall foreground objects at the first stage. Then, the proposals (Region of Interests, RoI) with high scores are sent to the second-stage detection heads for further refinement. One-stage approaches [122, 123] directly output each box and label in a per-pixel manner. In particular, with the help of focal loss and feature pyramid networks, one-stage approaches can even surpass the two-stage methods in several datasets. Both approaches need anchors for the regression of the bounding box. The anchors are the pixel locations where the objects are possibly emerging. Since two-stage approaches explicitly explore the foreground objects, where they have better recall than single-stage approaches. Two-stage approaches are more commonly used in open vocabulary object detection tasks.
@@
Instance segmentation to detect and segment each object, which goes beyond object detection. Most instance segmentation approaches focus on how to represent instance masks beyond object detection, which can be divided into two categories: top-down approaches [2, 124] and bottom-up approaches [125, 126]. The former extends the object detector with an extra mask head. The designs of mask heads are various, including FCN heads [2, 127], diverse mask encodings [128], and dynamic kernels [124, 129]. The latter performs instance clustering from semantic segmentation maps to form instance masks. The performance of top-down approaches is closely related to the choice of detector  [130], while bottom-up approaches depend on both semantic segmentation results and clustering methods [131]. Besides, there are also several approaches [132] using gird representation to learn instance masks directly.
@@
Query-based Object Detection and Segmentation. With the rise of vision transformers [3, 4], recent works mainly use transformer-based approaches in segmentation, detection, and video understanding. Compared with previous pixel-based approaches, transformer-based approaches have more advantages in cases of flexibility, simplicity, and uniformity [4, 133, 134, 135, 136, 137, 138, 139, 140, 49, 141]. One representative work is detection transformer (DETR) [4]. It contains a CNN backbone, a standard transformer encoder, and a standard transformer decoder. It also introduces the concepts of object query to replace the anchor design in pixel-based approaches. Object query is usually combined with bipartite matching [142] during training, uniquely assigning predictions with ground truth. This means each object query builds the one-to-one matching during training. Such matching is based on the matching cost between ground truth and predictions. The matching cost is defined as the distance between prediction and ground truth, including labels, boxes, and masks. By minimizing the cost with the Hungarian algorithm [142], each object query is assigned by its corresponding ground truth. For object detection, each object query is trained with classification and box regression loss [121]. For instance-wised segmentation, each object query is trained with classification loss and segmentation loss. The output masks are obtained via the inner product between object query and decoder features. Recently, mask transformers [143, 144, 145] further removed the box head for segmentation tasks. Max-Deeplab [143] is the first to remove the box head and design a pure-mask-based segmenter. It combines a CNN-transformer hybrid encoder [146] and a query-based decoder as an extra path. Max-Deeplab still needs many extra auxiliary loss functions. Latter, K-Net [144] uses mask pooling to group the mask features and designs a dynamic convolution to update the corresponding query. Meanwhile, MaskFormer [145] extends the original DETR by removing the box head and transferring the object query into the mask query via MLPs. It proves simple mask classification can work well enough for all three segmentation tasks. Then, Mask2Former [147] proposes masked cross-attention and replaces the cross-attention in MaskFormer. Masked cross-attention forces the object query only attends to the object area, guided by the mask outputs from previous stages. Mask2Former also adopts a stronger Deformable FPN backbone [148], stronger data augmentation [149], and multiscale mask decoding. In summary, query-based approaches are stronger and simpler. They still cannot directly detect novel classes but are widely used as base detectors and segmenters in open vocabulary settings.
@@
Large Scale Visual Language Pre-training. Better visual language pre-training can lead to a better understanding of semantics on given visual inputs [134, 150, 135]. Previous works [151, 152, 153] focus on cross-modality research via learning the visual and sentence-dense connection between different modalities. Some works [154, 153] utilize two-stream neural networks based on the vision transformer model, while several works [155, 156] adopt the single-stream neural network, where the text embeddings are frozen. The two-stream neural networks process visual and language information, respectively, and fuse them afterward by another transformer layer. Recently, most approaches [16, 17] adopt pure transformer-based visual language pre-training. Both CLIP [16] and Align [17] are concurrent works that explore extremely large-scale pre-training on image text pairs. In particular, with such large-scale training, CLIP demonstrates that the simple pre-training task of predicting which caption goes with which image can already lead to stronger generalizable models. These VLMs are the foundation of open vocabulary learning in different tasks, which means that the open vocabulary approaches aim to distill or utilize knowledge of VLMs into their corresponding tasks.
@@
where $e^{v}_{i}$ is the $i$-th vision embedding and $e^{t}_{j}$ is the $j$-th text embedding, and $p_{ij}$ is the prediction score of the $i$-th vision proposal predicted to the $j$-th class. The .,$\langle$.,$\rangle$ . , . ⟩ represents the dot product operation. The denominator is added with 1 because the background text embedding is set to all 0 by default. $C_{B}$ and $C_{N}$ are sets of base and novel classes, respectively. Finally, the class with the highest prediction score is chosen as the prediction. In practice, the vision embeddings are trained with base annotations to fit the text embeddings. Therefore, the model combines the knowledge of VLM and the learned visual features, and the detector/segmenter can detect/segment novel classes via semantically related text embeddings.
@@
HierKD [157]
OVOD
None
Bounding Boxes (Base)
CLIP-text
CLIP-vision + ATSS
Distills Global-level knowledge from the pre-trained CLIP model into the one-stage detector.
@@
Propose a wealy supervised approach that training rare classes with image-level annotations.
@@
Prompt-OVD [164]
OVOD
None
Bounding Boxes (Base)
CLIP-text
CLIP-vision + DETR
Propose RoI-based masked attention and RoI pruning techniques by utilizing CLIP visual features to improve the novel object classification.
@@
CLIP-vison + MaskFormer
@@
TCL [174]
OVSS
Captions
None
CLIP-text
CLIP-vision
Proposes a finer-grained contrastive loss for training without pixel-level annotations.
@@
Cat-Seg [179]
OVSS
None
Segmentation Masks (Base)
CLIP-text
CLIP-image
Jointly aggregates the image and text embeddings from CLIP to form the segmentation predictions.
SAN [180]
OVSS
None
Segmentation Masks (Base)
CLIP-text
CLIP-vision
Proposes a light-weight side adaptor network to extract the CLIP model’s knowledge.

### v4

@@ 2.3 Tasks, Datasets, and Metrics
$\bullet$ Datasets. For object detection, the common datasets are COCO [5] and LVIS [49]. Recently, a more challenging dataset, v3Det [50] with more than 10,000 categories, is proposed. For image segmentation, the most commonly used datasets are COCO [5], ADE20k [51], PASCAL-VOC 2012 [52], PASCAL-Context [53], and Cityscapes [54]. For video segmentation and tracking, the frequently used datasets are VSPW [55], Youtube-VIS [56], LV-VIS [57], MOSE [58], and TAO [59].
$\bullet$ Metrics. For detection tasks, the commonly used metrics are mean average precision (mAP) and mean average recall (mAP) for both base and novel classes. Among segmentation tasks, the commonly used metrics are mean intersection over union (mIoU) for semantic segmentation, mask-based mAP for instance segmentation, and panoptic quality (PQ) for panoptic segmentation.
@@ 2.4 Related Research Domains
$\bullet$ Open-Set Recognition. The concept of Open-Set Recognition (OSR) addresses the challenge of identifying unknown classes during classification tasks. In traditional classification systems, models are trained to recognize a finite set of known classes, but in real-world scenarios, they may encounter data that doesn’t belong to these predefined categories. During testing, OSR aims to classify known classes seen during the training and reject unknown classes that are unseen [31]. The main methods used for OSR can be broadly categorized into two types: discriminative models [34, 35, 36, 37, 60] and generative models [61, 62, 63, 64, 65, 66]. The discriminative models focus on differentiating between known classes and identifying unknown classes by enhancing the boundary or margin between these classes. The generative models are either instance generation-based [61, 62, 63, 64, 65] or non-instance generation-based [66]. They emphasize generating new instances or features to improve the ability of the system to recognize new, unseen classes during testing. Each category employs specific techniques and approaches to address the challenges inherent in open-set recognition, which involves dealing with classes not seen during the training phase.  [67] extends the OSR task to require further identifying novel classes.
$\bullet$ Open World Learning. Open world learning involves identifying and labeling new, unknown categories (novel unknowns) [38]. This process includes incrementally learning new categories as they are labeled and added to the system. The goal is to create a system that remains robust to unknown categories and adapts continually to include new information, balancing the risks associated with open spaces in the learning model. [38] first propose open word recognition. Recent works also explore the open world object detection [32, 68] task.
$\bullet$ Out-of-Distribution Detection. Out-of-distribution (OOD) detection methods can be structured into several categories [29]. Classification-Based Methods [69, 70, 71, 72]: These include output-based methods [69, 70], label space redesign [71], and OOD data generation techniques [72]. Density-Based Methods [73, 74, 75]: This category involves methods that detect OOD by modeling data density. Distance-Based Methods [76, 77]: These methods use distance metrics, typically in the feature space, to identify OOD instances. Reconstruction-Based Methods [78, 79, 80]: This approach achieves OOD detection that features reconstruction capabilities.
@@
$\bullet$ Zero-Shot Detection and Segmentation: This task aims to segment classes that have not been encountered during training. Two streams of work have emerged: discriminative methods [81, 82, 83, 84, 85] and generative methods [86, 87, 88, 89]. Representative works in this field include SPNet [81] and ZS3Net [87]. SPNet [81] maps each pixel to a semantic word embedding space and projects pixel features onto class probabilities using a fixed semantic word embedding [90, 91] projection matrix. On the other hand, ZS3Net [87] first trains a generative model to produce pixel-wise features for unseen classes based on word embeddings. With these synthetic features, the model can be trained in a supervised manner. Both of these works treat zero-shot detection and segmentation as a pixel-level zero-shot classification problem. However, this formulation is not robust for zero-shot learning, as text embeddings are typically used to describe objects/segments rather than individual pixels. Subsequent works [92, 88, 93, 86, 89, 94] follow this formulation to address different challenges in zero-shot learning. In a weaker assumption where unlabeled pixels from unseen classes are available in the training images, self-training [87] is commonly employed. Despite promising results, self-training often requires model retraining whenever a new class appears. ZSI [40] also employs region-level classification for bounding boxes but focuses on instance segmentation rather than semantic segmentation. More recently, PADing [86] proposes a unified framework to tackle zero-shot semantic segmentation, zero-shot instance segmentation, and zero-shot panoptic segmentation.
@@
Most approaches in open vocabulary learning are based on zero-shot learning settings, such as replacing the fixed classifier with language embeddings. However, these methods struggle to generalize well to novel classes due to the absence of novel class knowledge. Consequently, their performance is limited, and they are not practical for real-world applications.
$\bullet$ Long-tail Object Detection and Instance Segmentation. This task addresses the challenge of class imbalance in instance segmentation. Many approaches tackle this issue through techniques such as data re-sampling [49, 95], loss re-weighting [96], and decoupled training [97]. Specifically, some studies [95] employ image-level re-sampling. However, these methods tend to introduce bias in instance co-occurrence. To address this issue, other works [98] focus on more refined re-sampling techniques at the instance or feature level. Regarding loss re-weighting, most studies [96] rebalance the ratio of positive and negative samples during training. Additionally, decoupled training methods [97, 99] introduce different calibration frameworks to enhance classification results. Resolving long-tail object detection can lead to improved accuracy of rare classes. However, these methods currently cannot be applied to the detection of novel classes.
$\bullet$ Few-Shot Detection and Segmentation: Few-shot object detection aims to expand the detection capabilities of a model using only a few labeled samples. Several approaches [100, 101, 102, 21] have been proposed to advance this field. Notably, TFA [103] introduces a simple two-phase fine-tuning method, while DeFRCN [104] separates the training of RPN features and RoI classification. Moreover, SRR-FSD [105] combines multi-modal inputs while LVC [106] proposes a pipeline to generate additional examples for novel object detection and train a more robust detector. Few-shot segmentation comprises Few-Shot Semantic Segmentation (FSSS) and Few-Shot Instance Segmentation (FSIS). FSSS involves performing pixel-level classification on query images. Previous approaches [107, 108] typically build category prototypes from support images and segment the query image by computing the similarity distance between each prototype and query features. On the other hand, FSIS aims to detect and segment objects with only a few examples. FSIS methods can be categorized into single-branch and dual-branch methods. The former [109] primarily focuses on designing the classification head, while the latter [110, 111, 112] introduce an additional support branch to compute class prototypes or re-weighting vectors of support images. This support branch assists the segmenter in identifying target category features through feature aggregation. For example, Meta R-CNN [112] performs channel-wise multiplication on RoI features, and FGN [111] aggregates channel-wise features at three stages: RPN, detection head, and mask head. However, few-shot learning still requires examples of novel classes during training, and such data may not be available.
@@ 3 Methods: A Survey
Refer to captionViLD [12]BARON [113]LSeg [14]ZegFormer [114]MaskCLIP [48]OPSNet [115]OpenScene [116]ULIP [117]OVR-CNN [11]OpenSeg [46]CGG [47]DST-Det [118]PB-OVD [119]MaskCLIP+ [120]OVSeg [121]XPM [122]MAXI [123]PLA [124]GroupViT [125]SegCLIP [126]Mask-free OVIS [127]OpenSeeD [128]OpenSD [129]X-Decoder [130]FreeSeg [131]OVDiff [132]ODISE [133]Detic [134]OWLv2 [135]
Figure 4: Open vocabulary learning methods, organized by their tasks and approach types. We list several representative works here.
Figure 5: An illustration of a common architecture in open vocabulary object detection and segmentation. The vision model predicts a class embedding for each box/mask. The embeddings are compared to a set of class embeddings generated by a VLM text model like CLIP or ALIGN, using dot products. The class with the highest score is chosen as the predicted label for the object. Note that while humans define the set of possible object classes, the system only has access to a limited set of “base” classes during training.
Overview. In this section, we first review preliminary knowledge and vision language modeling by extending close-set detectors into open vocabulary detectors via VLMs in Sec. 3.1 and Sec. 3.2. Then we sequentially survey the six subsidiary tasks, including object detection  (Sec. 3.3), segmentation (Sec. 3.4), video understanding (Sec. 3.5), 3D scene understanding (Sec. 3.6), and closely related tasks (Sec. 3.7). Note that we only record and compare the most representative works. We list numerous works in Tab. I and Tab. IV. Moreover, since there are several similar tasks (Sec. 3.7), we also survey and compare other related tasks, including class agnostic detection and segmentation, open world object detection, and open-set panoptic segmentation.
Method taxonomy and relation of each subsection. We summarize the most commonly used methods among these different directions in Fig. 4. However, we cannot include all methods due to the wide range of tasks. In each subsection, we summarize different methods according to different problems, segmentation, detection, video understanding, and closely related topics. We cluster methods according to common practice, such as knowledge distillation and region text pre-training. We start with object detection since it was first proposed. For segmentation tasks, we ignore the specific setting in Sec. 3.4. We argue that despite these two directions, they may share similar ideas. However, most meta-architectures and tasks are different. Thus, we still review each direction individually. For video, 3D understanding, and other closely related topics, due to different task definitions and extra input cues, such as temporal information and multi-view inputs, we also survey two directions individually.
Motivation of survey organization. We argue that using the unified symbolic system to summarize all methods is hard. There are several reasons. (1), The task definitions and settings are different. In addition to different input and output formats, taking the open vocabulary segmentation as an example, there are at least three different settings [125, 47, 46] for open vocabulary segmentation. Thus, the design principles are different. For example, OpenSeg [46] uses supervised pretraining and contrastive losses while GroupViT [125] adopt unsupervised setting without mask annotations. (2), Different methods use different detectors or baseline models. Several works [11, 134] use R-CNN framework while several methods [47, 131] adopt query-based approaches. Several methods [136, 137] adopt the RPN’s features to develop their methods, while query-based methods cannot perform these designs. (3), Different methods may also use different datasets during training. Several methods [134, 138, 139] mainly explore the dataset effect on open-vocabulary detection or segmentation. Since the supervision signals are different, it will hard to put all methods in one unified symbolic system. In summary, our survey focus on more comprehensive review to extract the common features of various domains in open vocabulary learning.
@@ 3.1 Preliminary
Pixel-based Object Detection and Segmentation. In general, it can be divided into two aspects: semantic-level tasks and instance-level tasks, where the difference lies in whether to distinguish each instance. For the former, we take semantic segmentation as an example. It was typically approached as a dense pixel classification problem, as initially proposed by FCN [140]. Then, the following works are all based on the FCN framework. These methods can be divided into the following categories, including better encoder-decoder frameworks [141, 142], larger kernels [143], multiscale pooling [144, 145], multiscale feature fusion [146, 147, 148, 149], non-local modeling [150], and better boundary delineation [151, 152, 153, 154]. After the transformer was proposed, with the goal of global context modeling, several works propose the variants of self-attention operators to replace the CNN prediction heads [155, 156].
@@
For the latter, we take object detection and instance segmentation for illustration. Object detection aims to detect each instance box and classify each instance. It mainly has two different categories: two-stage approaches and one-stage approaches. Two-stage approaches [157, 2] rely on an extra region proposal network (RPN) to recall foreground objects at the first stage. Then, the proposals (Region of Interests, RoI) with high scores are sent to the second-stage detection heads for further refinement. One-stage approaches [158, 159] directly output each box and label in a per-pixel manner. In particular, with the help of focal loss and feature pyramid networks, one-stage approaches can surpass the two-stage methods in several datasets. Both approaches need anchors for the regression of the bounding box. The anchors are the pixel locations where the objects are possibly emerging. Since two-stage approaches explicitly explore the foreground objects, where they have better recall than single-stage approaches. Two-stage approaches are more commonly used in open vocabulary object detection tasks.
@@
Instance segmentation segmenting each object goes beyond object detection. Most instance segmentation approaches focus on how to represent instance masks beyond object detection, which can be divided into two categories: top-down approaches [2, 160] and bottom-up approaches [161, 162]. The former extends the object detector with an extra mask head. The designs of mask heads are various, including FCN heads [2, 163], diverse mask encodings [164], and dynamic kernels [160, 165]. The latter performs instance clustering from semantic segmentation maps to form instance masks. The performance of top-down approaches is closely related to the choice of detectors [166], while bottom-up approaches depend on both semantic segmentation results and clustering methods [167]. Besides, several approaches [168] use gird representation to learn instance masks directly.
@@
Query-based Object Detection and Segmentation. With the rise of vision transformers [3, 4], recent works mainly use transformer-based approaches in segmentation, detection, and video understanding. Compared with previous pixel-based approaches, transformer-based approaches have more advantages in cases of flexibility, simplicity, and uniformity [4, 169, 170, 171, 172, 173, 174, 175, 176, 58, 177, 178]. One representative work is detection transformer (DETR) [4]. It contains a CNN backbone, a standard transformer encoder, and a standard transformer decoder. It also introduces the concepts of object query to replace the anchor design in pixel-based approaches. Object query is usually combined with bipartite matching [179] during training, uniquely assigning predictions with ground truth. This means each object query builds the one-to-one matching during training. Such matching is based on the matching cost between ground truth and predictions. The matching cost is defined as the distance between prediction and ground truth, including labels, boxes, and masks. By minimizing the cost with the Hungarian algorithm [179], each object query is assigned by its corresponding ground truth. For object detection, each object query is trained with classification and box regression loss [157]. For instance-wised segmentation, each object query is trained with classification loss and segmentation loss. The output masks are obtained via the inner product between object query and decoder features. Recently, mask transformers [180, 181, 182] further removed the box head for segmentation tasks. Max-Deeplab [180] is the first to remove the box head and design a pure-mask-based segmenter. It combines a CNN-transformer hybrid encoder [183] and a query-based decoder as an extra path. Max-Deeplab still needs many auxiliary loss functions. Later, K-Net [181] uses mask pooling to group the mask features and designs a dynamic convolution to update the corresponding query. Meanwhile, MaskFormer [182] extends the original DETR by removing the box head and transferring the object query into the mask query via MLPs. It proves simple mask classification can work well enough for all three segmentation tasks. Then, Mask2Former [184] proposes masked cross-attention and replaces the cross-attention in MaskFormer. Masked cross-attention forces the object query only attends to the object area, guided by the mask outputs from previous stages. Mask2Former also adopts a stronger Deformable FPN backbone [185], stronger data augmentation [186], and multiscale mask decoding. In summary, query-based approaches are stronger and simpler. They cannot directly detect novel classes but are widely used as base detectors and segmenters in open vocabulary settings.
3.2 Vision Language Modeling
@@
Large Scale Visual Language Pre-training. Better visual language pre-training can lead to a better understanding of semantics on given visual inputs [170, 187, 171]. Previous works [188, 189, 190] focus on cross-modality research via learning the visual and sentence-dense connection between different modalities. Some works [191, 190] utilize two-stream neural networks based on the vision transformer model, while several works [192, 193] adopt the single-stream neural network, where the text embeddings are frozen. The two-stream neural networks process visual and language information and fuse them afterward by another transformer module. Recently, most approaches [15, 16] adopt pure transformer-based visual language pre-training. Both CLIP [15] and Align [16] are concurrent works that explore extremely large-scale pre-training on image text pairs. In particular, with such large-scale training, CLIP demonstrates that the simple pre-training task of predicting which caption goes with which image can already lead to stronger generalizable models. Several following works [194, 195, 196] aim to improve the CLIP training via mask image modeling, scaling up the training data and model size. These VLMs are the foundation of open vocabulary learning in different tasks, which means that the open vocabulary approaches aim to distill or utilize knowledge of VLMs into their corresponding tasks.
Moreover, in addition to achieving better zero-shot recognition, several works focus on designing better vision language models for language-related tasks, including visual question answering (VQA). Several works [197, 198, 199, 200, 201] explore how to better align caption loss and contrastive loss during the image-text pertaining. In particular, CoCa [198] adopts cross attention to connect caption generation part and contrastive learning. Recently, BLIP-2 [201] bootstraps vision-language pre-training from off-the-shelf frozen pre-trained image encoders and frozen large language models via a lightweight Querying Transformer.
$\bullet$ Visual Grounding Tasks. Visual grounding tasks aim to localize the specific objects according to given text descriptions. Referring segmentation tasks [171, 202, 203, 204] segment the specific object. Referring expression comprehension [205, 206, 207] localizes the bounding boxes of given object texts. Recent research is mainly based on two-stream networks: one for visual encoder [1, 3] and the other for text encoder [9]. These works focus on how to better match text features and visual features via different architectures. For referring segmentation, previous works [208, 209, 187] adopt “decoder-fusion”, where they design a separate decoder at the end of the two-stream networks. Recent works [210, 211] explore the “encoder-fusion” to directly fuse the text features into a visual backbone before mask prediction. For referring expression comprehension, similar approaches [205, 212] are adopted such as encoder-fusion. Recently, several works [213, 214] aim to unify visual grounding and object detection. In particular, Grounding DINO [214] unifies detection and grounding in one framework. It uses stronger detectors with multiple dataset pre-training, which achieves strong results for many downstream tasks. Compared with open vocabulary learning, visual grounding tasks require visual text matching with specific text descriptions. Open vocabulary learning tasks require the model to automatically detect, segment and recognize new objects without the given text information, such as class names, which is more challenging.
@@
where $e^{v}_{i}$ is the $i$-th vision embedding and $e^{t}_{j}$ is the $j$-th text embedding, and $p_{ij}$ is the prediction score of the $i$-th vision proposal predicted to the $j$-th class. The .,$\langle$.,$\rangle$ . , . ⟩ represents the dot product operation. The denominator is added with one because the background text embedding is set to all 0 or make it learnable. $C_{B}$ and $C_{N}$ are base and novel classes, respectively. Finally, the class with the highest prediction score is chosen as the prediction. In practice, the vision embeddings are trained with base annotations to fit the text embeddings. Therefore, the model combines the knowledge of VLM and the learned visual features, and the detector/segmenter can detect/segment novel classes via semantically related text embeddings.
@@
Propose a weakly supervised approach that is training rare classes with image-level annotations.
@@
PADing [86]
OVSS
None
Segmentation Masks (Base)
CLIP-text
Mask2Former
Introduce a generative model to synthesize features for unseen categories and achieve universal open-vocabulary segmentation.
@@
CLIP-vision + MaskFormer
@@
TABLE II: Feature summarization of current open vocabulary object detection approaches.
Method Common Features
Knowledge Distillation Distil the aligned visual-text knowledge of VLM into object detection. Focus on distilling loss design and targe proposal generation.
Region Text Pre-training Use large-scale and easily available text-image pairs to pre-train a better and universal detector and finetune on the target datasets
Training with More Balanced Data Adopt extra data or balanced datasets as an augmentation to train the classification head.
Prompting Modeling Adopt different prompting modeling to better transfer the VLM knowledge into detector.
Region Text Alignment Design a better region-text alignment method to better align the region visual features (RoIs) into text features of VLM.